OpenStudy (anonymous):

Find an equation of the line. slope 3; through (-1,7) write the equation in standard form.

OpenStudy (texaschic101):

y = mx + b slope(m) = 3 (-1,7) x = -1 and y = 7 now sub and solve for b and tell me what you get

OpenStudy (anonymous):

7=-3+b 10=b

OpenStudy (texaschic101):

correct. So your equation in slope intercept form is y = 3x + 10 but you need it in Ax + By = C y = 3x + 10 -- subtract 3x from both sides -3x + y = 10 -- multiply by -1 to make x positive 3x - y = -10 Do you understand ?

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Yes I do , thank you very much !

OpenStudy (texaschic101):

no problem :)